# Avoiding Catastrophe in Continuous Spaces by Asking for Help

### v2

@@ C.2 Proof roadmap
    Bound the total number of buckets and therefore the total number of queries (Lemma C.1).
@@
    Bound the suboptimality on a single time step based on the bucket length and the Lipschitz constant (Lemma C.2).
@@
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma C.4), with Lemma C.3 as an intermediate step. This captures the total amount of suboptimality.
@@
    The bound from Lemma C.4 is in additive form, whereas the objective is multiplicative. Lemma B.7 factors the objective into one term for the optimal payoffs and one term for the suboptimality in the form of Lemma C.4.
@@
    Combine Lemmas C.2, C.4 and B.7 to bound the total regret (Lemma C.5).
@@
    Prove Theorem 5.2 using Lemma C.1, Lemma C.5, and some light arithmetic.
C.3 Proof
Lemma C.1.
Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries.
@@ C.3 Proof
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))+2$.
@@
Therefore the total number of buckets ever in existence is $(2+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)+2\leq(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$, so Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries. ∎
Lemma C.2.
For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})$.
@@
If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$. Applying Lemma B.5 completes the proof. ∎
Lemma C.3.
If $\pi^{*}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
@@
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$. Then there exists $s^{\prime}\in B$ (as defined in Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $t\in V_{T}$, we have $\pi^{*}(s_{t})\neq a_{t}=\pi^{*}(s^{\prime})$. Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$. Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$. Then $\max(S_{B})<\max(B)\leq\max(S_{B}^{\prime})$, so $S_{B}\neq S_{B^{\prime}}$. Thus the number of segments in $\pi^{*}$ is at least $\mathcal{B}_{V}^{\prime}$. Since the number of segments in $\pi^{*}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$. ∎
Lemma C.4.
@@
Lemma C.5.
For any $\mathbf{s}\in S^{T}$, any Lipschitz $\mu$, any $\pi^{*}$ with at most $K$ segments, and any $g:\mathbb{N}\to\mathbb{N}$, Algorithm 2 satisfies
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq\frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)$
@@
Let $\mu_{0}=4L/g(T)$. Suppose there exists $t\in[T]$ such that $\mu^{*}(s_{t})<\mu_{0}$. Then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\mu_{0}-0\leq\frac{6L}{g(T)}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2L\operatorname*{len}(B_{t})$ for $t\in V_{T}$. Then by Lemma C.2,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Note that $\operatorname*{len}(B_{t})\leq 1/g(T)$ for all $t\in[T]$, so $c_{t}\in[0,\mu_{0}/2]$. Thus by Lemma B.7,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-3L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
Then by Lemma C.4,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\frac{6LKT}{g(T)^{2}}$
and so
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\frac{6LKT}{g(T)^{2}}$
@@
See 5.2
@@
Lemma C.1 implies the query bound. Since $KT\geq T\geq g(T)$, Lemma C.5 implies that for any such $\mathbf{s},\mu$, and $\pi^{*}$,
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq$ $\ \frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)=\frac{6L}{g(T)}\cdot\frac{TK}{g(T)}$
$=$ $\ \frac{6LTK}{T^{2c}}=6LKT^{1-2c}$

### v3

@@ D.3 Proof roadmap
    Bound the total number of buckets and therefore the total number of queries (Lemma D.1).
@@
    Bound the suboptimality on a single time step based on the bucket length and $L$ (Lemma D.2).
@@
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the total amount of suboptimality.
@@
    As in the proof of Theorem 5.2, Lemma B.3 transforms the multiplicative objective into an additive form. Lemma D.5 bounds the additive objective using Lemmas D.2 and D.4.
@@
    Combining Lemmas D.5 and B.3 bounds the regret (Lemma D.6).
@@
    Theorem D.2 directly follows from Lemmas D.1 and D.6.
D.4 Proof
Lemma D.1.
Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
@@ D.4 Proof
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{diam}(\mathbf{s})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\mathbf{s})+2$.
@@
Therefore the total number of buckets ever in existence is $(\operatorname*{diam}(\mathbf{s})+2)g(T)+2\leq(\operatorname*{diam}(\mathbf{s})+4)g(T)$, so Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries. ∎
Lemma D.2.
For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$.
@@
If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$. Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by the local generalization assumption,
$\mu(s_{t},a_{t})=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$
as required. ∎
Lemma D.3.
If $\pi^{m}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
@@
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$. Then there exists $s^{\prime}\in B$ (as defined in Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$. Since $t\in V_{T}$, we have $\pi^{m}(s_{t})\neq a_{t}=\pi^{m}(s^{\prime})$. Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$. Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$. Then $\max(S_{B})<\max(B)\leq\max(S_{B^{\prime}})$, so $S_{B}\neq S_{B^{\prime}}$. Thus the number of segments in $\pi^{m}$ is at least $\mathcal{B}_{V}^{\prime}$. Since the number of segments in $\pi^{m}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$. ∎
Lemma D.4.
@@
Lemma D.5.
Under the conditions of Theorem D.2, Algorithm 4 satisfies
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
@@
For $t\not\in V_{T}$ we have $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ by definition, and Lemma D.2 implies that $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$. Thus
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
Then by Lemma D.4,
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
@@
Lemma D.6.
Under the conditions of Theorem D.2, Algorithm 4 satisfies $R_{T}\leq\frac{2LKT}{g(T)^{2}}$.
@@
Let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
$R_{T}=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Applying Lemma D.5 completes the proof. ∎
Theorem D.2 follows from Lemma D.1 and Lemma D.6.
Appendix E Other proofs
Proposition E.1.
Suppose $\mu$ is $L$-Lipschitz continuous and $\pi^{m}$ is optimal, i.e., $\mu(s,\pi^{m}(s))=\max_{a\in A}\mu(s,a)$ for all $s\in S$. Then $\pi^{m}$ satisfies local generalization with constant $2L$.
Proof.
For any $s,s^{\prime}\in S$, we have
$\mu(s,\pi^{m}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{m}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
$\geq$ $\ \mu(s^{\prime},\pi^{m}(s))-L||s-s^{\prime}||$ ($\pi^{m}$ is optimal for $s^{\prime}$)
$\geq$ $\ \mu(s,\pi^{m}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
$=$ $\ \mu^{m}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{m}(s)$)
Since $\pi^{m}$ is optimal for $s$, we have
$\mu^{m}(s)+2L||s-s^{\prime}||\geq\mu^{m}(s)\geq\mu(s,\pi^{m}(s^{\prime}))$
Thus $-2L||s-s^{\prime}||\leq\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)\leq 2L||s-s^{\prime}||$. This is equivalent to $|\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)|\leq 2L||s-s^{\prime}||$, completing the proof.